Solve the following problems. The activation energy for the reaction below equals 1.0 x 105 J/mol. Given k 2.5 x 103 sec1 at 332 K, find k at 375 K. 1. N20s (g)-2NO (g) + 0a(g) Based on information in problem 1, find the temperature at which k is twice as large as it is at 332K. 2. Please show all your work and write neatly please! Thank you guys! ??
Answer #1

1) ln (K1 / K2) =Ea / R [1/T2 - 1/T1]
K1= 2.5*10-3 sec-1 ; T1 = 332 K
K2 = ? ; T2 = 375 K
Ea = 1*105 J/mol ; R = 8.314 J/mol K
ln (2.5*10-3 / K2) =1*105 / 8.314 [1/375 - 1/332]
ln (2.5*10-3 / K2) = -4.154216
2.5*10-3 / K2 = e-4.154216 = 0.015698
K2 = 0.15925 sec-1

2) T1 = 332 K ; K1 = 0.15925 sec-1
T2 = ? ; K2 = 2*0.15925 = 0.3185 sec-1
use the same formula
ln (0.15925 / 0.3185) = 1*105 / 8.314 [1/T2 - 1/332]
-5.7628*10-5 = 1/T2 - 1/332
T2 = 338.4759 ~ 338.5 K